If Q2 and Q3 are separated by a distance of 20 cm and have the same charge of 50μC, what is the magnitude of the force between them?


1
Expert's answer
2022-03-20T18:51:29-0400

The Coulomb's law says

"F=k\frac{q_1q_2}{r^2}"

"F=9\times 10^9\times \frac{50\times 10^{-6}\times 50\times 10^{-6}}{0.20^2}=562.5\:\rm N"